OpenStudy (anonymous):

Evaluate log2 1 over 64.

OpenStudy (anonymous):

−6 −32 6 32

OpenStudy (whpalmer4):

Can you evaluate \[\log_2 64\]?

OpenStudy (anonymous):

I'm not sure how, I'm willing to learn though :)

OpenStudy (whpalmer4):

remember that \[\log_b n = a\]where \[b^a=n\]

OpenStudy (whpalmer4):

so \[\log_2 64 = a \text{ such that } 2^a = 64\]

OpenStudy (whpalmer4):

2*2=4 2*2*2=8 2*2*2*2=16 2*2*2*2*2=32 2*2*2*2*2*2=64 \[\log_2 64 = \]?

OpenStudy (anonymous):

2^6?

OpenStudy (whpalmer4):

well, 2^6= 64, yes. but a = 6

OpenStudy (***[isuru]***):

hate to interrupt guys... i know whpalmer is the senior one.. but u guys r missing something... what Brad tryingto say is \[\log_{2} \frac{ 1 }{ 64 }\]

OpenStudy (whpalmer4):

No, we're not missing that. it's easier to explain the positive log case first...

OpenStudy (whpalmer4):

but thanks for making sure!

OpenStudy (***[isuru]***):

u can use \[\log \frac{ a }{ b } = \log a - \log b\]

OpenStudy (***[isuru]***):

WOW u r awesome then !!

OpenStudy (whpalmer4):

such a wise response certainly deserves a medal :-)

OpenStudy (whpalmer4):

@Brad1996 do you see why the \[\log_2 64 = 6\]?

OpenStudy (anonymous):

Log still confuses me... I mean what does it mean when you ask the log of something?

OpenStudy (whpalmer4):

Okay, good question. Logs are like exponents, except with a fixed choice for the base. As you recall, with an exponential like \(a^b\) \(a\) is the base, and \(b\) is the exponent. When we decide to take logarithms, we choose 1 value for the base, that's the base of the logarithm, in this problem it happens to be 2. Now, finding the log of \(n\) is just finding the exponent that you need to raise your base to to get \(n\): \[\log_b a = n\]\[b^n = a\] \[\log_2 64 = 6\]\[2^6 = 64\]\[\log_2 1 = 0\]\[2^0 = 1\] We can't ever have the log of a negative number, or 0, because there is no exponent to which we could raise any base to give us either of those answers.

OpenStudy (anonymous):

thanks, still a bit confusing but I'm sure I'll get it through practice :)

OpenStudy (anonymous):

I'm going to post another question and try to work it through, could you help me through it?

OpenStudy (whpalmer4):

One of the nifty things about logs is that it turns multiplication and division problems into addition and subtraction problems, and raising numbers to powers turns into multiplication. Wait! You haven't finished here yet.

OpenStudy (whpalmer4):

We did the log of 64, but we need the log of 1/64 to answer this question! Do you remember the property of exponents that \[a^{-n} = \frac{1}{a^n}\]

OpenStudy (anonymous):

Yeah I think so

OpenStudy (anonymous):

so it would be -6?

OpenStudy (anonymous):

because 2^-6 = 1/64?

OpenStudy (whpalmer4):

Yep! Excellent!

OpenStudy (anonymous):

Sweet :D

OpenStudy (whpalmer4):

The other guy made an important point about logs that we could have used here, would you like me to elaborate on it? I'll still look at your other problem, of course.

OpenStudy (anonymous):

sure

OpenStudy (whpalmer4):

Logs turn multiplication and division into addition and subtraction. For example, to compute 32*8, we could find \(\log_2 32\), \(\log_2 8\), add them, and take the anti-log \(2^n\). What are those two logs?

OpenStudy (anonymous):

you lost me there bro lol

OpenStudy (whpalmer4):

Well, just figure out the two logs for me: \[\log_2 32 = \]\[\log_2 8 = \]

OpenStudy (anonymous):

2^3 = 8 2^5 = 32

OpenStudy (anonymous):

OS being a bit slow for me

OpenStudy (whpalmer4):

Yeah, me too! :( Okay, so if \(2^3=8\), what is \(\log_2 8 =\) If \(2^5 = 32\), what is \(\log_2 32 = \)